OpenStudy (anonymous):

find and equation that has x= -2, x=5, and x= -1 as solutions

OpenStudy (jdoe0001):

so the solutions are x = -2 => (x+2) = 0 x = 5 => (x-5) = 0 x = -1 => (x+1) = 0 so those were the factors of the original polynomial or (x+2)(x-5)(x+1) = 0 to get the equation in standard form, multiply them over

OpenStudy (anonymous):

multiply them over?

OpenStudy (jdoe0001):

well, multiply them :)

OpenStudy (jdoe0001):

3 roots/solutions, means 3rd degree polynomial

OpenStudy (anonymous):

yep. im lost

OpenStudy (jdoe0001):

\(\bf (x+2)(x-5)(x+1) = 0 \implies (x+2)\times(x-5)\times(x+1) = 0 \)

OpenStudy (anonymous):

so thats the equation?

OpenStudy (anonymous):

i can get to x^2 - 5x + 2x - 10 (x+1) = 0

OpenStudy (jdoe0001):

\(\bf (x+2)(x-5)(x+1) = 0 \implies (x^2-5x+2x-10)(x+1)=0\\\quad \\ (x^2-3x-10)(x+1)=0 \implies (x^3-3x^2-10x)+(x^2-3x-10)\)

OpenStudy (jdoe0001):

\(\bf (x^2-3x-10)(x+1)=0 \implies (x^3-3x^2-10x)+(x^2-3x-10)\\\quad \\ x^3-3x^2+x^2-10x-3x-10\)

OpenStudy (jdoe0001):

then just add up like-terms

OpenStudy (anonymous):

thanks!!

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

i ended up with \[x ^{3}-2x ^{2}-13x-10=0\]